OpenStudy (bizzeizzy):

what is the equation of the line perpendicular to the line y=2/3x+2 that passes through the point (1, 5)?

OpenStudy (bizzeizzy):

\[y=\frac{ 2 }{ 3 }x+2\]

jimthompson5910 (jim_thompson5910):

what is the slope of that given line?

OpenStudy (bizzeizzy):

\[\frac{ 2 }{ 3 }\]

jimthompson5910 (jim_thompson5910):

flip the fraction: 2/3 ----> 3/2 flip the sign: 3/2 ----> -3/2

jimthompson5910 (jim_thompson5910):

the perpendicular slope is -3/2

jimthompson5910 (jim_thompson5910):

the perpendicular line goes through (1,5) (x,y) = (1,5) x = 1 y = 5

jimthompson5910 (jim_thompson5910):

so we have this info m = -3/2 x = 1 y = 5 you plug those values into y = mx+b and solve for b

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (lunymoony):

b = 6?

OpenStudy (lunymoony):

i dont think this is right

OpenStudy (lunymoony):

but Im out of my head tonight so Im not sure

OpenStudy (bizzeizzy):

so y-5=-3/2 (x-1)

OpenStudy (bizzeizzy):

right?

jimthompson5910 (jim_thompson5910):

now solve for y

OpenStudy (bizzeizzy):

y=1/2(13-3x)

jimthompson5910 (jim_thompson5910):

that's another way of expressing \[\Large y = -\frac{3}{2}x + \frac{13}{2}\]

jimthompson5910 (jim_thompson5910):

what I wrote is in slope intercept form m = -3/2 is the slope b = 13/2 is the y-intercept

OpenStudy (bizzeizzy):

ohh okay thank you for the help!

jimthompson5910 (jim_thompson5910):

you're welcome